Tom and Jerry have to sort and organize all the nails on their father's workbench. Tom can sort 220 nails in 2 hours. Jerry can sort 150 nails in 3 hours. How long will it take them to sort 500 nails if they work together (round to the nearest hundredth)?
Answer by ankor@dixie-net.com(22740)

Tom and Jerry have to sort and organize all the nails on their father's workbench.
Tom can sort 220 nails in 2 hours. Jerry can sort 150 nails in 3 hours.
How long will it take them to sort 500 nails if they work together (round to the nearest hundredth)?
:
Let t = no. of hrs for 500 nails working together
:
Find how many nails each can sort in an hour
Tom: 220/2 = 110 per hr
Jerry: 150/3 = 50 " "
:
110t + 50t = 500
160t = 500
t = 500/160
t = 3.125 hrs which is 3 + .125(60) = 3 hrs 7.5 minutes
:
:
We can confirm this
3.125(110) + 3.125(50) = 500